OpenStudy (anonymous):

2cos(x)+radical2=0 solve the following for radians

OpenStudy (anonymous):

start with \[\cos(x)=-\frac{\sqrt{2}}{2}\] and solve that

OpenStudy (anonymous):

there are an infinite number of answers, but the first one you will run across is \(x=\frac{3\pi}{4}\) next will be \(\frac{5\pi}{4}\)

OpenStudy (anonymous):

take the inverse cosin=-radical2/2?

OpenStudy (anonymous):

take the inverse cosin=-radical2/2?

OpenStudy (anonymous):

you can do that if necessary, but that will only give you one answer easier just to visualize it. take a look at the unit circle on the last page of the attached cheat sheet. cosine is the first coordinate, so find the places on the unit circle where the first coordinate is \(-\frac{\sqrt{2}}{2}\)

OpenStudy (anonymous):

thank you!